OpenStudy (anonymous):

I NEED HELPPP- WILL GIVE MEDAL What is the resulting ordered pair if the value of the independent variable is -1/2? f(x)=12x^2-8x

OpenStudy (anonymous):

f(x)=12x^2-8x f(-1/2)=12(-1/2)^2 - 8(-1/2) f(-1/2)=12(1/4) -8(-1/2) f(-1/2)=12/4 + 8/2 f(-1/2)=3 +4 f(-1/2)=7 so pair (-1/2, 7)

OpenStudy (johnweldon1993):

The dependent variable here...is 'y' ...because What 'y' equals...is dependent on what 'x' equals So the independent variable is 'x' ...and we are told that the value of the independent variable is (-1/2) so we would plug in (-1/2) for 'x' in our equation \[\large 12(-\frac{ 1 }{ 2 })^2 - 8(-\frac{ 1 }{ 2 })\] And solve for 'y'